OpenStudy (anonymous):

A medal to anyone who answers this question (: Evaluate the expression below when x = 2 3x^2-2/x+4 A) 1 B) 5/4 C) 5/3 D) 17/3

OpenStudy (anonymous):

Which of the following is your equation and you should be specific next time \[3x^{2}-\frac{2}{x+4}=11\frac{2}{3}\] or \[3x^{2-\frac{2}{x+4}}=62.9\]

OpenStudy (anonymous):

It's 3x^2 -2 over x+4

OpenStudy (anonymous):

Then the answer is 1